### v3

# Automated Static Vulnerability Detection via a Holistic Neuro-symbolic Approach

Abstract.
In this paper, we present MoCQ, a novel neuro-symbolic framework that combines the complementary strengths of Large Language Model (LLM) and classic vulnerability checkers to enable scalable, automated vulnerability detection. The key insight is to leverage an LLM to automatically generate vulnerability patterns and translate them into detection queries. Specifically, MoCQ incorporates an iterative loop in which an LLM refines queries based on carefully designed feedback information. The resulting queries are then executed to analyze large codebases and detect vulnerabilities. We evaluated MoCQ on 12 vulnerability types across four programming languages. MoCQ achieved comparable precision and recall compared to expert-developed queries, with significantly less expert time needed. MoCQ also uncovered 46 new vulnerability patterns that experts missed, each representing an overlooked vulnerability class. MoCQ further discovered seven previously unknown vulnerabilities in real-world applications.
Neuro-symbolic Analysis; Static Vulnerability Detection
journalyear: 2025ccs: Security and privacy Software and application security
1. Introduction
Pattern-based static analysis (i.e., vulnerability checker) identifies vulnerabilities by matching source code against known vulnerable patterns, without requiring program execution. Compared to dynamic techniques (Pham et al., 2017; Trickel et al., 2023; Lee et al., 2021), it scalably analyzes large codebases with high code coverage, and requires less setup effort (e.g., no need for software deployment). It has been widely adopted in practice (Backes et al., 2017; Yamaguchi et al., 2014; Shi et al., 2024; Shcherbakov et al., 2023) and has discovered numerous high-impact vulnerabilities (e.g., Log4Shell (Services, 2021; MITRE Corporation, 2021)). For instance, CodeQL (CodeQL, 2024a) is extensively integrated into GitHub’s CI/CD pipelines and is employed to analyze large-scale applications such as the Chrome browser (Project, 2025). Joern (Yamaguchi et al., 2014; Backes et al., 2017) has been recognized as one of the most widely used tools for automated vulnerability discovery (Shiri Harzevili et al., 2024).
Despite its success, the effectiveness of this approach hinges entirely on the quality of vulnerability patterns. These patterns describe vulnerable code structures or behaviors and are used to match new vulnerability instances in target codebases. However, the creation of these patterns is largely manual and demands substantial human effort. This process is not only time-consuming (e.g., several weeks or more (Fröberg, 2023)), but also prone to inaccuracies due to limited human knowledge. This significantly reduces the effectiveness of detection (N/A, 2024; cod, 2024). As codebases evolve and new threats emerge, keeping detection patterns accurate and up to date requires ongoing maintenance and engineering effort.
To alleviate this burden, researchers have alternatively explored learning-based approaches that analyze source code without execution. Most deep learning (DL)-based methods operate at the function level to classify functions as vulnerable or not. However, they often lack interprocedural or project-wide context (Li et al., 2021b; Russell et al., 2018), and require a large volume of labeled vulnerabilities for training (Chakraborty et al., 2021), which may not be available. More recently, several works (Zhou et al., 2024; Lu et al., 2024; Ding et al., 2025) have used Large Language Models (LLMs) to directly analyze code without relying on traditional program analysis. Leveraging massive training corpora, LLMs can recognize typical (vulnerable) usage patterns and reason about code semantics across diverse programming contexts. While promising, LLMs face challenges such as token length limits, inconsistent reasoning, and susceptibility to hallucinations, which limit their reliability in analyzing large, complex software.
The distinct challenges of classic and learning-based methods suggest that a hybrid neuro-symbolic approach [^1] Neuro refers to learning-based analysis (particularly LLMs in this paper) leveraging neural models for semantic reasoning; symbolic refers to classic static analysis using symbolic representations and formal logic systems. is a promising direction. While recent works have taken steps toward such integration, they fall short of the level of automation. For example, IRIS (Li et al., 2025) and Artemis (Ji et al., 2025) use LLMs to extract taint sources and sinks, and incorporate them into existing detection patterns to enhance the coverage of vulnerability searches. However, these approaches only extend some parts of base patterns developed by humans, which still require manual effort and are prone to human error.
In this paper, we design the first static, automated neuro-symbolic vulnerability detection framework, called MoCQ (Model-Generated Code Queries), that combines the complementary strengths of LLMs and classic static analysis. The core idea is to leverage LLMs to automatically generate end-to-end vulnerability patterns and incorporate the patterns into static program analysis. Such a neuro-symbolic design not only preserves the scalability and rigorousness of classic static analysis, but also gets rid of the time-consuming pattern engineering.
However, LLM-driven generation of vulnerability patterns faces multiple challenges. LLMs often do not have enough knowledge of domain-specific languages (DSLs) used by static vulnerability detection tools (e.g., Scala-like DSL for Joern (Yamaguchi et al., 2014) and SQL-like DSL for CodeQL), resulting in uncompilable queries. Specifically, our study shows that pretrained, commercial LLMs, e.g., GPT-4o (OpenAI, 2024), often fail to generate valid queries, frequently producing syntax errors or triggering execution-time exceptions. To deal with this challenge, MoCQ adopts a technique, specifically DSL subsetting, to refine DSL by selecting a core set of features to guide the automated query generation, without sacrificing expressiveness. The intuition is that such a refined subset can be better understood by pretrained LLMs while still preserving the core semantics due to the redundant nature of the DSL.
Even when the LLM-generated queries are syntactically correct and executable, they may fail to capture vulnerabilities effectively. Specifically, they may trigger semantic errors (i.e., fail to report vulnerabilities), be overly specific (i.e., overfitting to known instances via matching exact variable names or fixed control flow paths), or be overly general (i.e., underfitting by matching many irrelevant code cases). To deal with these, MoCQ adopts a feedback-driven approach, which incorporates fine-grained feedback from a symbolic query validator for iterative query refinement. The query validator executes the LLM-generated queries and monitors the exhibited behaviors, including syntax errors, execution exceptions, and semantic errors. Specifically, we instrument the query execution runtime to obtain a block-level program state of the query execution to precisely locate the errors and inconsistencies. These are then used for the LLM to refine the incorrect query. The validator further uses heuristics to detect if a query is too specific and instructs the LLM to generalize it; it also instructs LLMs to eliminate false positives (if any) to improve the precision.
We implemented MoCQ for two state-of-the-art static detection tools, Joern (Yamaguchi et al., 2014) and CodeQL (CodeQL, 2024a). MoCQ is language-agnostic and can generate patterns for all programming languages supported by these tools. We extensively evaluated MoCQ on 12 representative vulnerability types (e.g., use-after-free and cross-site scripting) on four popular programming languages. Our results show that, given only a few (e.g., 5-18) vulnerability examples, MoCQ can efficiently generate detection queries within hours. These queries demonstrate strong detection capability comparable to that developed by security experts. MoCQ-generated queries successfully detected seven new vulnerabilities in real-world applications.
Beyond detection performance, MoCQ uncovered 46 new vulnerability patterns missed by expert-developed queries, with five patterns already integrated into upstream tool repositories. Each missing pattern often results in overlooking entire vulnerability classes. Our further analysis revealed complementary strengths of LLMs and human experts: LLMs excel at capturing subtle linguistic nuances that experts consistently overlook, while experts demonstrate superior high-level reasoning for complex vulnerabilities.
This paper makes the following contributions.
  • We proposed a novel way to split the tasks for neural and symbolic components to achieve their complementary benefits.
  • We developed a DSL subsetting technique and a feedback-driven approach in MoCQ, which automatically generates end-to-end vulnerability patterns.
  • We demonstrated that MoCQ could automatically generate vulnerability patterns for real-world vulnerabilities and achieve comparable results with patterns from security experts.
  • MoCQ uncovered 46 new patterns and seven new vulnerabilities.
2. Background and Motivation
2.1. Pattern-based Static Analysis
Pattern-based static analysis is a representative and practical approach among various static analysis techniques. Tools (e.g., Joern (Yamaguchi et al., 2014; Backes et al., 2017) and CodeQL (CodeQL, 2024a)) express vulnerability patterns as structured queries written in tools’ domain-specific languages (DSLs). They use language-specific frontends to translate source code into graph-like program representations such as control-flow graphs (CFG) and code property graphs (CPG) (Yamaguchi et al., 2014). Security analysts then write precise queries that operate on these representations to identify vulnerabilities. Joern models programs with CPGs (joe, 2025), while CodeQL relies on code databases. The DSLs differ as well: Joern employs a Scala-like language, whereas CodeQL uses SQL-like syntax. Static analyzers enforce strict syntax and semantic checks during query execution to ensure accuracy and soundness.
Example: A Prototype Pollution Query. JavaScript is a prototype-based language where objects inherit properties and methods through a mechanism known as prototypes. Modifying the prototype of one object can affect all other objects that inherit from the same prototype, resulting in prototype pollution. For example, an attacker can access the prototype via p=obj["__proto__"] and modify with it using p["toString"]="Hacked". As a result, any object inheriting from the fundamental prototype Object.prototype now has tampered toString method. This type of vulnerability can lead to arbitrary code execution, privilege escalation, and denial of service (Shcherbakov et al., 2023; Li et al., 2021a; Cornelissen et al., 2024; Liu et al., 2024; Kang et al., 2022).
  1 // find obj["__proto__"] in p=obj["__proto__"]  2 def objProto = cpg.call  3   .where(_.name(Operators.assignment)).argument(2).isCall.arrayAccess   4   5 // find p["toString"] in p["toString"]="Hacked"  6 def objProp = cpg.call  7   .where(_.name(Operators.assignment)).argument(1).isCall.arrayAccess   8   9 // find a data-flow path from objProp to objProto  10 objProp.array.reachableBy(objProto).filter(...)
Figure 1. Example real-world query in Joern, simplified for clarity.
Figure 1 shows a simplified version of a real-world Joern query designed to detect JavaScript prototype pollution (Fröberg, 2023). [^2] The original query is more complex, including checks for user-controlled values, property filtering, etc. It comprises three main steps: (1) identifying the object prototype (lines 1-3), (2) identifying the property access (lines 5-7), and (3) verifying whether there is a data connection between them (lines 9-10). For example, step (1) uses Joern’s dot operator (".") to chain operations, where each stage filters results produced by the previous one. It begins by selecting assignment expressions via Operators.assignment, [^3] In Joern, assignment is modeled as a form of call operation. extracts the right-hand side via argument(2), and narrows results to array-style accesses using arrayAccess().
Query Analysis. Writing high-quality patterns is non-trivial. For example, the query in Figure 1 took about three weeks to draft and an additional month to refine (que, 2023). Experts may miss attack variants or write overly general patterns, causing false negatives and false positives (N/A, 2024; cod, 2024). For instance, attackers can bypass the detection by combining property access and assignment in one expression, like obj["proto"]["toString"]="Hacked". This also happens in the official CodeQL queries maintained by GitHub Security.
This work thus explores an automated approach to generating vulnerability patterns using LLMs. When a new threat is publicly disclosed, a developer can rapidly generate comprehensive vulnerability patterns to scan their codebases, without the long delays traditionally required for manual pattern engineering.
2.2. Challenges
While using LLMs to generate queries may seem intuitive, our preliminary study uncovers substantial challenges. We tasked GPT-4o (OpenAI, 2024) with generating vulnerability queries for five types of vulnerabilities using three prompting strategies: zero-shot, few-shot, and chain-of-thought (CoT). In each case, we provided an input vulnerability example as the detection goal. Out of 300 cases, only one successful query was produced in the few-shot setting to retrieve the input vulnerability example. We omit full results due to space constraints. We summarize four key technical challenges. Below, we show a LLM-generated query for JavaScript prototype pollution that failed to execute in Joern’s analysis engine.
An invalid query generated in few-shot prompting:   1 ...   2 val assign = code.call // omitting ‘cpg‘   3   .nameExact("assignment") // using wrong operator name
DSL Syntax Correctness. The LLM-generated queries frequently raise grammar and syntax errors, which stem from the model’s incomplete understanding of the DSL grammar. Specifically, LLMs typically lack sufficient exposure to the syntax, semantics, and typical usage of complex DSLs. For example, Joern queries based on the code property graph usually require the root object cpg. However, the query generated above did not follow this requirement. This issue is further aggravated by the hallucinations of LLMs.
Runtime Execution Correctness. Queries must execute within the domain-specific runtime (i.e., analysis engine) and interact meaningfully with structured program representations. LLM-generated queries often use incorrect or non-existent function calls, missing parameters, and incompatible data types or operations, resulting in execution-time exceptions. For example, the nameExact operation in the above query used "assignment" instead of "<operator>.assignment", which led to execution-time exceptions after we first manually fixed the syntax error.
Semantic Validity. Even when queries are syntactically correct and executable, ensuring their semantic validity remains challenging. They must satisfy runtime constraints and capture the correct traversal logic. Executable queries may still fail to reflect intended program behaviors, leading to semantic errors, where the query runs without crashing but misses the vulnerable code.
Precision-recall Balance. Similar to how security analysts develop queries, LLM-generated queries must be not only correct but also comprehensive and precise. Patterns that are too specific (i.e., overfitting) may miss variants of the same vulnerability, resulting in high false negatives, whereas overly general patterns (i.e., underfitting) can produce excessive false positives. Achieving this balance between precision and recall remains a fundamental challenge.
3. MoCQ
We design MoCQ, a novel static neuro-symbolic system. Its workflow is outlined in Figure 2. MoCQ first analyzes open knowledge such as documentation and tool implementations to extract the query DSL, and constructs a core DSL subset (§ 3.1). MoCQ then generates vulnerability queries (§ 3.2) through a feedback loop, with a trace-driven symbolic query validator to refine (§ 3.3) and optimize (§ 3.4) the queries. Finally, MoCQ outputs the generated query, which can be applied to detect new vulnerabilities. We summarize two key techniques that solve the aforementioned challenges.
Query DSL Extraction and Subsetting. To help generate syntactically- and semantically-correct queries, we first obtain the formal specifications of the query DSLs used by the static analysis tools, which also define the desired output structure for LLMs. MoCQ extracts the DSL specifications—including DSL grammars, data types, compatible APIs, and their functionalities—by automatically parsing the online documentation and implementations. However, directly using all extracted DSL specifications would overwhelm the LLM because of the DSL’s complexity, and would still not enable the LLM to efficiently generate valid queries. We thus propose a language subsetting technique, which selects a core subset of DSL features from the extracted full set. This is based on our observation that query DSL often contains redundant features, for example, those that can be equally expressed by other language features. Such a DSL subset could be better understood by the LLMs to significantly reduce the query generation complexity.
Iterative Feedback Loop. Atop the core DSL subset, MoCQ employs an LLM to generate queries and iteratively refines them based on fine-grained feedback from our symbolic query validator. Specifically, we instrument query execution runtime, which executes and analyzes the generated queries to locate the exact locations or causes of the syntax and semantic errors. Beyond that, the validator also tracks the intermediate program states on the query execution traces and validates whether the generated query could retrieve the given vulnerability examples. During such an iterative process, MoCQ is able to gradually generate valid queries that can analyze complex real-world programs. Moreover, the validator incorporates heuristics to detect potential overfitting in queries and guides the LLM to generalize them. It also instruments the LLM to refine the query to eliminate the false positives.
Figure 2. The workflow of MoCQ.
3.1. DSL Specification Extraction
To help LLMs generate syntactically and semantically valid queries, MoCQ extracts formal specifications of the query DSLs used by static analysis tools. Rather than using few-shot learning (which provides a few query examples) or fine-tuning (which requires a large amount of labeled data), we guide the model using a concise DSL specification. This specification includes both the grammar, which defines the structural rules of the query language, and the semantics, which describe the meaning and behavior of APIs, data types, and parameters. This extraction is performed once per static analysis tool.
3.1.1. DSL Extraction
The extraction process targets both syntax and semantics. To this end, MoCQ first analyzes the static analysis tool’s documentation (e.g., online use instructions) and source code (e.g., API implementations and comments). Documentation typically contains structured content (e.g., HTML tables) that describes grammar, operations, and usage patterns. We first crawl the online documentation of the tools and prompt an LLM to extract grammar rules, data types, operations, and API usage. For example, Joern’s website (N/A, 2025f) lists the query operations (termed “Steps”) along with their descriptions.
Online documentation only briefly mentions the APIs’ behaviors. We seek more precise and comprehensive definitions from their code implementations, which offer ground-truth definitions through function signatures, type annotations, inline comments, etc. Based on the results of document analysis (e.g., keywords, API name), we prompt an LLM to retrieve relevant source code files from each tool’s implementation, then to automatically summarize the functionality of the related functions.
3.1.2. DSL Subsetting
The extracted query DSL specification is complex and could overwhelm the model. For example, static analysis tools like Joern and CodeQL have thousands of APIs. Each API can have multiple parameter choices, resulting in an exponential number of combinations of these rules and operations. Even though MoCQ has known the precise description for each operation, our evaluation (details in § 5.5.6) demonstrates that providing LLMs with the full-set DSL does not efficiently produce good queries.
We design a language subsetting technique to resolve the DSL complexity issue. We observe that the DSL specification contains significant redundancy not strictly necessary for constructing useful queries. A small, expressive subset is often sufficient for most vulnerability detection tasks. Instead of exposing the entire DSL, we restrict query generation to this smaller, more tractable subset, which reduces complexity and improves accuracy. Concretely, the subsetting process operates at the granularity of APIs, which can strike a balance between expressiveness and control—it preserves flexibility for query construction while keeping the DSL surface concise and understandable. MoCQ first associates typical parameter choices with each API. It then analyzes the full set of DSL APIs and prompts an LLM to decide and select these important, necessary DSL APIs for common usage. This is feasible since the LLM can refer to features’ intended behaviors or functionalities to evaluate if one is necessary.
We also provide a small collection of queries to help MoCQ learn the core DSL subset. Here, the LLM only needs to base on them to select the API subset from the full set, which is much simpler than generating new queries using few-shot prompting with query examples. Nevertheless, if a security analyst identifies additional needs or corner cases, they can perform a one-time augmentation to expand the DSL subset.
3.1.3. DSL Prompting
We provide the core DSL subset to the LLM to guide query generation. We adopt a grammar prompting technique (Wang et al., 2023b) to concisely express the grammar. Specifically, we supply an abstract grammar specification of the DSL using the Backus-Naur Form (BNF) (McCracken and Reilly, 2003), which is a standard way of defining language syntax. Although BNF is itself domain-specific, it appears more frequently in training data than custom query DSLs. Moreover, the BNF structure is simpler and more concise than real query examples, while still conveying rich information about DSL usage. We also specify the subset semantics like APIs to the LLM in a compact form, including their signature, data types, and API functionalities.
Figure 3 presents a simplified grammar for Joern using BNF. It details the fundamental grammar rules, such as the cpg_start that initializes a traversal query, and how various components connect to form complete expressions. The grammar rules explicitly define the structure and permissible compositions of the DSL, guiding models to produce syntactically valid statements. For example, by following the grammar, an LLM knows to begin queries with cpg followed by appropriate method chaining through step_chain elements. The grammar also clarifies how to construct complex queries using operations like filter_step and complex_step, enabling more sophisticated code analysis patterns.
{grammar}
¡traversal¿ ::= ¡cpg_start¿ ”.” ¡step_chain¿
¡cpg_start¿ ::= ”cpg” — ”cpg.method” — ”cpg.call” — …
¡step_chain¿ ::= ¡step¿ — ¡step¿ ”.” ¡step_chain¿
¡step¿ ::= ¡filter_step¿ — ¡complex_step¿
¡filter_step¿ ::= ”.where(” ¡predicate¿ ”)”
¡complex_step¿ ::= ”.filter(” ¡predicate¿ ”)” — ”.reachableBy(” ¡traversal¿ ”)” — …
¡predicate¿ ::= ”_.name(” ¡reference¿ ”)” — ¡identifier¿
¡reference¿ ::= ”Operators.” ¡operator¿
¡operator¿ ::= ”assignment” — ”arrayAccess” — ”fieldAccess” — …
Figure 3. Simplified BNF grammar for Joern, extracted by MoCQ.
3.2. Query Generation
Atop the core DSL subset, MoCQ generates detection queries as described in Algorithm 1. Each time, MoCQ is tasked to generate a query ($Q^{t}$) for a specific vulnerability type ($t$). It takes as input (1) a task description ($D^{t}$) that contains a natural language explanation of the vulnerability type, and (2) a small set of vulnerability examples in the type ($E^{t}=\{e^{t}_{i}\mid i=1,\ldots,m\}$). To make query generation more manageable, MoCQ first produces a per-example query ($Q^{t}_{i}$) that could analyze and retrieve the corresponding example $e^{t}_{i}$. Then, MoCQ performs optimizations to address overfitting and underfitting (line 8). Finally, the system merges all per-example queries into a unified per-type query ($Q^{t}=\bigcup_{i=1}^{m}Q^{t}_{i}$) for effective real-world vulnerability detection (line 14).
The vulnerability examples play a crucial role in our design. First, they help LLMs understand the semantics and patterns associated with each vulnerability type, especially for unseen threats. Second, MoCQ uses these examples as test cases to validate the generated queries. We believe this is a reasonable and practical assumption for deploying MoCQ in real-world scenarios because vulnerability examples are often available, for example, from the CVE database and patches. To this end, MoCQ takes a program slice for each vulnerability example—retaining the necessary code syntax and semantics for query generation.
As an initial step, we include the general workflow of the vulnerability type in the task description. Instead of composing a single monolithic query with hundreds or thousands of operations, MoCQ applies a chain-of-thought prompting strategy to decompose the detection task into smaller, well-defined subtasks. For example, to detect prototype pollution, MoCQ splits the task into four steps: (1) identifying object property modifications, (2) determining if user-controlled input affects the property name, (3) inspecting the property name itself, and (4) checking for sanitization. Importantly, MoCQ can automatically perform this decomposition based on the task descriptions, without manual intervention.
1
Input : Task-Description $D^{t}$, Vulnerability examples $E^{t}$
Output : Query $Q^{t}$
2 $QList$leftarrow [ ]
3 foreach $e^{t}_{i}\in E^{t}$ do
4  $PState\leftarrow null$
5  for $i\leftarrow 1$ to $MaxN$ do
6     $Q^{t}_{i}$leftarrow LLMQueryGen ($e^{t}_{i}$, $D^{t}$, PState)
7     error, PState $\leftarrow$ Validation ($Q^{t}_{i}$, $e^{t}_{i}$)
8     if not error then
9        $Q^{t}_{i}$leftarrow Optimize ($Q^{t}_{i}$)
10        QList.add ($Q^{t}_{i}$)
11        break
12     end if
13
14  end for
15
16 end foreach
[…]
18 return $Q^{t}$
19
20
21 function Validation (Q, e):
22  $S\leftarrow null$
23  foreach $B\in Q$ do
24     S $\leftarrow$ Execute (B, e, S) // Capture fine-grained runtime info.
25
26  end foreach
27 return S.err, S.State
28 end
Algorithm 1 Iterative query generation in a feedback loop.
3.3. Iterative Query Refinement
Given the difficulty of query generation, it is almost impossible to produce a qualified query in a single LLM attempt. Therefore, MoCQ incorporates an iterative feedback loop (shown as the for loop on lines 4-11 of Algorithm 1) to debug and refine the query using fine-grained feedback from a trace-driven symbolic query validator. This process terminates either when a valid query is generated or when the maximum attempt threshold is reached.
3.3.1. Symbolic Validation
MoCQ first asks the LLM to reflect on its past query output, suggest an improved version, and then retry the task (Renze and Guven, 2024; Li et al., 2023). This helps mitigate randomness and hallucinations. Different from prior self-reflection techniques, MoCQ employs a trace-driven symbolic query validation to provide fine-grained feedback information to the LLM.
In particular, the generated query ($Q^{t}_{i}$) can be represented as a sequence with $n$ blocks denoted as $[B_{1},B_{2},\ldots,B_{n}]$. The query produces a trace when executing. When applying the $Q^{t}_{i}$ to analyze the application containing the vulnerability example $e^{t}_{i}$, MoCQ records the intermediate program states for its blocks along the trace. Specifically, when the first $j$ blocks have been executed, MoCQ collects the set of in-scope variables and their runtime values by instrumenting the query execution runtime. The corresponding program state is formulated as $S_{j}=\{(var,val_{j})\mid var\in B_{\leq j}\}$, where $var$ represents the in-scope variable till block $B_{j}$, and $val_{j}$ is the corresponding value at current time. To help locate the root cause of failure, we provide the LLM fine-grained program state information with annotated trace values $PState=[S_{1},S_{2},\ldots,S_{n}]$. Our design of feedback further includes checks for syntax errors, execution exceptions, and semantics errors.
Syntax Validation. Syntax validation checks if the query complies with the extracted DSL grammar. The static analysis tool itself provides a runtime for the queries and is naturally equipped with grammar validation capability. However, it is often coarse-grained without pinpointing the exact error locations. Therefore, we instrument and hook the grammar validator to obtain the corresponding grammar rule of the errors. Alternatively, one can leverage off-the-shelf grammar parser (e.g., constructing an ANTLR (Parr, 2013)).
Execution Validation. Similarly, MoCQ refers to the collected DSL specifications for execution exceptions. There are multiple types of semantic errors such as undefined variables, missing attributes, type mismatches, or incorrect API usage. Our symbolic validator checks these behaviors and outputs the error locations. Besides merely identifying the errors, MoCQ also searches on the collected DSL semantics to suggest (potentially correct) operations by fuzzily matching the name similarities. For example, it could suggest "<operator>.assignment" as a similar argument for "assignment" to fix the exactName exception mentioned in § 2.2.
Semantic Validation. MoCQ also evaluates and inspects the semantic behaviors of the query execution. Specifically, MoCQ debugs the runtime behavior of the query with the account of the vulnerability example $e^{t}_{i}$. This leverages the knowledge of the vulnerability example to provide feedback to the LLM for refining the query. A straightforward approach is to check whether the query as a whole retrieves the vulnerability example (i.e., boolean feedback). However, this is insufficient for helping an LLM pinpoint the root cause when the query fails to retrieve the example. MoCQ thus filters in only the program states corresponding to the vulnerability example from the whole-project analysis, enabling the LLM to cross-check query behavior and self-correct potential issues.
3.4. Query Optimization
We develop three optimizations to mitigate query underfitting and overfitting and to improve query execution efficiency.
3.4.1. Eliminating FPs for Precision
When using the query $Q^{t}_{i}$ to analyze the project containing the input vulnerability example $e^{t}_{i}$, it might return false positives. A false positive is often introduced at a certain code block $B_{j}$. MoCQ thus similarly leverages the program state $PState$ to help the LLM reason about how the false positive is introduced and propose fixes.
In our evaluation, we take a best-effort manner to manually collect all true-positive vulnerabilities in a project as the example set $E^{t}$, therefore, all cases outside are considered false positives. In practice, non-vulnerable code is much more prevalent than vulnerable code, so a report from an underdevelopment query is more like to be a false positive. Recent research has shown it is possible to leverage an LLM to assess if a case is false positive or not (Li et al., 2025). We leave this as a future work.
3.4.2. Generalization for Recall
To mitigate the overfitting issue, MoCQ generalizes the queries to not just focus on a specific example. This is especially important to make the queries capable of real-world detection later on. As an initial step, we design a few heuristics to assess if a query is overfitting. First, MoCQ checks whether the query relies on exact constant values (e.g., hardcoded variable names and string literals) that are unique to a specific example. Such reliance limits the applicability of the query to other contexts. Second, MoCQ examines whether the query includes structural patterns or constraints that are overly specific to the layout of a particular program, such as deep AST chains or unique call sequences. Once any such situation is found, MoCQ employs the LLM to abstract queries by asking it to express the query in a more general form to resolve the identified overfitting situation. Such generalization improves the reusability of queries and enables the detection of vulnerability variants.
3.4.3. Merging for Efficiency
After generating individual queries for different vulnerability examples, MoCQ performs query merging to consolidate detection queries and improve efficiency. A naive approach would simply apply a logical OR operation to combine all per-example queries—sequentially running each one independently. However, this can lead to performance degradation, as many queries share redundant and repetitive conditions. To address that, MoCQ introduces an LLM-assisted query merging stage. All individual queries $Q^{t}_{i}$ are given to an LLM, which attempts to merge them into a single optimized query ($Q^{t}$) while preserving their detection capability. The merged query is passed through the symbolic query validator for all vulnerability examples to verify its correctness and effectiveness (this process is omitted from Algorithm 1). If the merged query fails validation on any of the examples, MoCQ further undergoes a similar feedback loop to iteratively improve the merging process.
4. Implementation
We implemented our approach in a highly modular way and integrated it with Joern (Yamaguchi et al., 2014) and CodeQL (CodeQL, 2024a). The hardware requirements for running MoCQ align with the standard recommendations for Joern (Joern, 2024) and CodeQL (CodeQL, 2024b), typically 16 GB RAM and 4 to 8 CPU cores.
We realized the runtime program state tracking by instrumenting the query execution runtime. In particular, we enhanced the query parsing process to locate syntax errors. For the other two types, we hooked the internal language interpreter. Specifically, these query languages’ interpreters typically contain a main interpretation loop that switches over the instruction types and invokes specific handlers. We thus added additional hookers for selected handlers to collect variables and their runtime values.
5. Evaluation
We extensively evaluate MoCQ to answer the following questions.
  • RQ1: New Vulnerabilities and Patterns. How effective is MoCQ in discovering new vulnerabilities in real-world applications using new vulnerability patterns?
  • RQ2: Query Effectiveness. How effective are the MoCQ-generated queries for detecting vulnerabilities? How do they compare to related approaches?
  • RQ3: Performance and Cost. What are the performance and cost of MoCQ?
  • RQ4: Ablation Study. How does each component of MoCQ contribute to its performance?
5.1. Experimental Setup
Dataset. MoCQ is applicable to a wide range of vulnerability types and programming languages. Our evaluation focuses on representative vulnerabilities across four popular programming languages (C, Java, PHP, and JavaScript), spanning systems, enterprise, and web domains. This selection covers distinct vulnerability classes from memory safety issues to injection attacks, such as use-after-free (UAF), overflow, injection, etc. We select these in part because expert-developed queries are available, enabling direct and meaningful comparisons. For each vulnerability type, we selected projects that contain documented vulnerabilities for query generation and testing. We collected from public sources (including the CVE database, GitHub issues, and patch commits) all known vulnerabilities to facilitate evaluation of both true positives and false negatives. In total, our dataset includes 306306306 vulnerabilities. The projects have an average size of  [^14] K lines of code and are thus well suited to assess the scalability of MoCQ.
Since vulnerabilities within the same project may share similarities, using them across query generation and evaluation could lead to data leakage. We thus partition the dataset at the project level into two isolated sets with manual verification to prevent data leakage: D1 includes projects containing 212212212 vulnerabilities for query generation; D2 include projects containing 949494 vulnerabilities for effectiveness evaluation; to assess generalizability, we introduce D3, consisting of 21 latest version projects not used in D1 or D2.
Configuration of MoCQ. MoCQ is a complex system with multiple configurable components, and we will systematically investigate their impact in § 5.5. Unless otherwise noted, our default setting involves generating queries from vulnerability examples drawn from ten randomly selected projects in D1 (146 out of 212 vulnerabilities), using the Claude 3.7 Sonnet model and Joern; the model is set with a temperature of 0.2 and a maximum token size of 10,00010,00010 , 000. We ran the experiments three times with an attempt threshold of 1,0001,0001 , 000 iterations on a CPU server equipped with dual 48-thread Intel Xeon processors and 187 GB RAM.
Table 1. New vulnerabilities discovered by MoCQ.
Project Vul. Type Experts’ MoCQ Status
forkcms (N/A, 2025d) PHP SQLi Reported
SyliusCmsPlugin (N/A, 2025a) PHP XSS $\times$ Reported
parse-static-imports (N/A, 2025e) JS Cmd. $\times$ Acknowledged
es2015-proxy (N/A, 2025c) JS Proto. $\times$ Acknowledged
web-worker (N/A, 2025g) JS Proto. Reported
koa-send (N/A, 2025h) JS Proto. Acknowledged
content-type (N/A, 2025b) JS Proto. $\times$ Reported
5.2. RQ1: New Vulnerabilities and Patterns
We apply the queries generated under the default setting on the latest project dataset (D3). MoCQ-generated queries effectively discover seven zero-day vulnerabilities, shown in Table 1. For comparison, we collected existing expert-developed queries from the official tool (joe, 2025) as well as from recent publications (Li and Meng, 2021; Fröberg, 2023). On the same dataset, these queries identified three vulnerabilities (which were also detected by MoCQ-generated queries) and missed the other four. The ability to outperform expert-developed queries in new vulnerability detection demonstrated the high usability of MoCQ. We responsibly reported the vulnerabilities and to date, three have been acknowledged.
Beyond new vulnerabilities, MoCQ discovered 46 new vulnerability patterns, including both vulnerability-specific patterns and three general patterns applicable across entire programming languages, as shown in Table 2. We define an atomic operation—such as a specific API call, expression type, or data flow—that is not recognized by expert-developed queries as a new pattern, and construct a minimal PoC vulnerability example to validate each new pattern. For example, four new PHP type juggling mechanisms were found by MoCQ, including built-in functions array_search and array_flip, where implicit type casting could occur; a new general way to propagate JavaScript data flows via Reflect.set, which could benefit detection of different vulnerability types.
Although expert-developed queries for these fundamental vulnerability types had undergone multiple rounds of refinement (see § 5.4), MoCQ still uncovered critical patterns that had been overlooked. Each pattern represents a blind spot in the detection logic, and missing even one can result in an entire class of vulnerabilities going undetected. This highlights the challenge of manually designing comprehensive query coverage, even for well-known bug types. In contrast, MoCQ benefits from LLMs’ ability to capture and enumerate subtle vulnerability behaviors. We submitted the new patterns to the query maintainers, and to date, five have been merged into upstream tools’ repositories.
Table 2. Overview of 46 new vulnerability patterns uncovered by MoCQ, including 34 distinct operations in the row of phar*.
Vul. Type Pattern Functionality Status
General PHP htmlspecialchars Ctx-sensitive sanitization Reported
PHP type. in_array Implicit type casting Merged
PHP type. array_search Loose comparison Merged
PHP type. array_flip Implicit type casting Reported
PHP type. case in switch Type coercion in case matching Reported
PHP deser. *phar Object deserialization Reported
PHP deser. copy Object deserialization Reported
General JS Object.assign Data propagation Reported
General JS Reflect.set Data manipulation Merged
JS proto. Object.assign Data propagation Merged
JS proto. Object.defineProperty New property manipulation Reported
JS proto. obj.__defineGetter__ New property manipulation Merged
C/C++ UAF Reallocation via malloc Data propagation Reported
Table 3. Vulnerability detection using Joern on the testing dataset.
MoCQ Experts’ MoCQ+Experts’ Pure LLM Generation Time Analysis Time
Vul. Type TPtotal TP FP Rec. Prec. TP FP Rec. Prec. TP FP Rec. Prec. TP FP Rec. Prec. Exp.’ MoCQ Exp.’ MoCQ
PHP SQLi 19 16 14 0.84 0.53 17 21 0.89 0.45 18 23 0.95 0.44 9 45 0.47 0.17 2d 15.3h 23m 33m
[…]
PHP type. 6 4 22 0.67 0.15 3 18 0.50 0.14 4 24 0.67 0.14 2 23 0.33 0.08 7d 5.2h 14m 20m
[…]
JS proto. 11 10 8 0.91 0.56 7 12 0.64 0.37 11 13 1.00 0.46 4 39 0.36 0.09 7w 21.4h 19m 26m
[…]
JS XSS 11 8 15 0.73 0.35 9 11 0.82 0.45 11 12 1.00 0.48 6 34 0.55 0.15 - 10.5h 20m 27m
[…]
C/C++ heap ov. 4 3 6 0.75 0.33 3 2 0.75 0.60 3 5 0.75 0.38 1 18 0.25 0.05 4w 9.4h 11m 14m
[…]
Java SQLi 3 3 5 1.00 0.38 2 2 0.67 0.50 3 3 1.00 0.50 1 15 0.33 0.06 3w 8.5h 12m 16m
[…]
Overall 94 73 119 0.79 0.39 65 111 0.67 0.40 81 137 0.85 0.39 45 341 0.44 0.11 - 160.1h 183m 260m
5.3. RQ2: Detection Effectiveness
We evaluate MoCQ on D2 and compare to related approaches.
5.3.1. Detection Results of MoCQ
D2 contains true vulnerabilities and per-type breakdown is shown as $TP_{total}$ in Table 3. Overall, MoCQ-generated queries could detect a broad range of vulnerabilities across PHP, JavaScript, C/C++, and Java ecosystems. Out of 94 known vulnerabilities, MoCQ-generated queries successfully detected 73 cases with 119 false positives. We also computed the recall ($\frac{TP}{TP_{total}}$) and precision ($\frac{TP}{TP+FP}$). MoCQ achieved a strong overall recall of 0.79 and a precision of 0.39. The precision, especially the false positive rate, is reasonable in practice, particularly for statically analyzing dynamic languages like PHP (Luo et al., 2022) and JavaScript (Shcherbakov et al., 2023) and unmanaged languages like C/C++ (Sui and Xue, 2016). As we will show in § 5.3.2, expert-developed queries have similar precision.
We analyzed the 21 vulnerabilities missed by MoCQ-generated queries. These false negatives stem from multiple root causes. There are 14 cases caused by incomplete patterns, including missing taint source specifications—particularly involving customized third-party sources—missing vulnerable operations, etc., which could have been detected if a more comprehensive pattern had been generated. Besides, four cases involved incomplete data flows where the targets of dynamic function calls were missing from the project’s program representations. This occurred because the static construction of representations failed to properly handle dynamic language features and indirect calls. Three cases were caused by flaws in type-related issues. For instance, inaccurate type information led to one false negative in PHP type juggling and two in JavaScript prototype pollution. The latter two categories could not be fixed by refining the patterns.
5.3.2. Comparison to Expert-developed Queries
We evaluated the collected expert-developed queries on the same dataset to thoroughly assess the effectiveness. Overall, MoCQ-generated queries achieved comparable performance to expert-developed queries, despite the latter requiring substantial security expertise and engineering effort. As shown in Table 3, expert-developed queries detected slightly fewer true positives (65 vs. 73) and produced slightly fewer false positives (111 vs. 119), resulting in an average recall of 0.67 and a precision of 0.40. A detailed comparison reveals that MoCQ and expert-developed queries detected 54 vulnerabilities in common, indicating a substantial overlap in their detection capabilities. In seven vulnerability types (PHP XSS, PHP type juggling, JavaScript command injection, C/C++ UAF, C/C++ heap overflow, Java SQLi, Java API misuse), MoCQ-generated queries detected all vulnerabilities found by the expert-developed queries.
However, our in-depth manual analysis of the queries reveals the different cognitive strengths of LLMs and human experts. In particular, LLMs excel at capturing subtle low-level linguistic nuances that humans might overlook, while human experts are better at high-level conceptual reasoning and understanding broader vulnerability contexts. For example, expert-developed queries all overlooked PHP’s case-insensitivity (e.g., echo is the same as eCHo), which was detected by MoCQ’s strength in syntactic nuances. Conversely, expert queries showed superior design in identifying complex multi-step SQL injection (Dahse and Holz, 2014) that require understanding broader application context.
This observation motivated us to investigate whether combining both sets of queries could lead to improved detection. We applied Joern using the union of MoCQ-generated and expert-developed queries, resulting in a combined approach that achieved a high recall of 0.85 and a precision of 0.39 on average. These results highlight that while MoCQ alone may not always outperform expert-developed queries, combining them yields a robust and highly effective detection strategy. Following that, we conducted a preliminary experiment to apply MoCQ to refine the expert-developed query for JavaScript prototype pollution as an example. MoCQ ran for around one hour and automatically revised the query, which in return detected 3 previously missed vulnerabilities. This highlights that MoCQ can not only automatically generate detection queries but also improve an existing expert-developed query.
5.3.3. Comparison to Other Approaches
We compare MoCQ to other related approaches on D2. In particular, we include a diverse set of methods, including a pure LLM-based approach, a prior neuro-symbolic method, and state-of-the-art static methods.
Pure LLM-based Detection. This approach directly provides the source code to an LLM and asks it to assess if there is any vulnerability. Unlike MoCQ, there is no LLM-based tool that can perform whole project analysis for large codebases, to the best of our knowledge. We thus take the common practice by separately checking individual code snippets (source code file) (Ding et al., 2025; Mohajer et al., 2024) with Claude 3.7 Sonnet. To handle non-determinism across multiple LLM runs, we assign the final label based on majority agreement—if at least two out of three runs agree, we adopt that label for the case.
The evaluation results in the last part of Table 3 show lower precision and recall. This is primarily due to the lack of cross-file information, such as global variables and inter-procedural function calls defined outside the current file. This leads to many false positives and false negatives. For instance, detecting JavaScript prototype pollution requires tracking data flow from external user input to object property manipulation. It is thus often infeasible when analyzing a single file in isolation. Providing more context to the LLM may potentially improve the results, however, LLMs inherently have the scalability issue for large-scale codebases.
IRIS. IRIS (Li et al., 2025) is a related work that combines LLM-identified taint specifications with existing queries. We consider its contribution orthogonal or complementary to that of MoCQ. IRIS is not directly applicable to the evaluated languages or vulnerability types for an end-to-end comparison. We thus incorporated its main technique and evaluated how it could help resolve our false negatives. The results showed that IRIS could successfully identify missing data sources for three false negative cases. Nevertheless, IRIS does not have a holistic, automated way of synthesizing end-to-end patterns.
Other State-of-the-art Static Methods. Our comparison to expert-developed queries demonstrates MoCQ’s advantage of automation. We are also interested in how MoCQ performs compared to other general, state-of-the-art, non-query-based static analysis tools. However, these tools often specialize in single vulnerability types, unlike MoCQ’s broad multi-language, multi-vulnerability support. It is non-trivial and time-consuming to set up each specialized tool for all covered vulnerability types. We thus select three representative, state-of-the-art analyzers (LChecker (Li and Meng, 2021) for PHP type juggling, Silent-Spring (Li et al., 2022) for JavaScript prototype pollution, and Cred (Yan et al., 2018) for C/C++ UAF) and evaluated them on D2. These tools employ different design and implementation choices (e.g., LLVM-based, CFG-based). We thus believe they effectively represent the diversity and sophistication of modern static analysis paradigms. Unlike MoCQ, all these approaches require a lot of engineering of vulnerability patterns.
LChecker reported six true positives and 37 false positives, while MoCQ reported four true positives and 22 false positives. LChecker benefits from a more precise type inference algorithm than the built-in type information available in Joern, which helps improve recall. However, in dynamic languages like PHP, many variable types often cannot be resolved statically. As a result, LChecker conservatively assigns multiple possible types, which can lead to over-approximation and more false positives. A more robust type inference would help Joern and MoCQ improve both precision and recall. For prototype pollution detection, Silent-Spring reported ten true positives and 12 false positives, achieving the same recall as MoCQ (91%) but a lower precision of 45%. For UAF, Cred reported four true positives with fewer false positives (four). This is because Cred leverages LLVM-based pointer analysis, which enables more accurate resolution of memory access patterns in C/C++ programs than Joern’s graph-based analysis. Nevertheless, MoCQ represents a significant step in automating pattern-based analysis.
5.4. RQ3: Performance and Cost
We measure the performance and cost of MoCQ, and present the results in the last four columns of Table 3.
Query Generation Time. As shown in Table 3, MoCQ required only a few hours to generate queries for each vulnerability type, significantly shorter than previous best practices that relied heavily on expert engineering. Generally, types with more input examples took longer to process. On average, it took 0.9 hours to construct a query on a single vulnerability example across different types.
We analyzed the Git commit history in the open-source repositories of the expert-developed queries. We find that experts often use multiple commits to develop and revise the queries, and the commit history often spans days to a few weeks. We acknowledge that Git commit history does not precisely reflect the actual, continuous engineering effort involved. For instance, the time span between the first and last commit may overestimate or underestimate the true effort, as substantial work may have occurred before the first commit or between commits without being recorded. Nevertheless, it serves as a reasonable proxy for comparing development timelines. As a comparison, MoCQ, as an automated solution, can run continuously and shorten the amount of time to construct queries.
Analysis Time. As shown in the last two columns of Table 3, MoCQ’s automated queries took 260 minutes in total to complete analysis on the tested projects, moderately longer than the 183 minutes required by expert-developed queries. This difference arises because most MoCQ patterns are broader and thus take more time to execute. Analysis time depends on application complexity and query design, but the overall duration of MoCQ’s remains practical for real-world use.
API Cost. Query generation per vulnerability type under the default setting (Claude Sonnet 3.7) typically takes less than 150 iterations with costs \$36 USD on average for LLM inference. Once generated, these queries can scan unlimited codebases without additional API cost.
5.5. RQ4: Ablation Study
We conduct a comprehensive ablation study to measure the individual contributions of the techniques on the testing dataset. We use the F1 score ($2\cdot\frac{Prec.\cdot Rec.}{Prec.+Rec.}$) as our evaluation metric, as it captures both precision and recall, and is well-suited for imbalanced data where non-vulnerable code is much more than vulnerable code. Due to space constraints, we summarize the main results and omit those for query merging and model temperature, though both of which were evaluated.
5.5.1. Vulnerability Examples
To understand how many vulnerability examples are needed for MoCQ to produce good detection results, starting from vulnerability examples in two projects, we gradually include more projects as inputs for MoCQ to generate queries. We then apply the queries to the testing dataset and compute the F1 scores. This incremental setup helps us understand the impact of additional projects on the detection effectiveness. We conduct our experiments at the project granularity rather than the vulnerability granularity due to similar data leakage concerns (discussed in § 5.1).
Our results show that F1 scores generally increase as the number of input projects grows, though the stabilization point varies by vulnerability type. For instance, JavaScript XSS queries reach stable high F1 scores with around four projects, while JavaScript prototype pollution requires closer to ten projects to stabilize. Overall, using about ten projects as input is sufficient for MoCQ to generate high-quality queries across most vulnerability types. This corresponds to roughly 5-18 vulnerability examples, depending on the type.
5.5.2. FP Elimination
During the iterative query generation, MoCQ treats those cases outside the generation dataset as false positives and uses them to refine the queries (§ 3.4.1). To understand the impact of such false positive signals, we experimented by disabling them as a comparison and showed the results as $W/O_{FP}$ in Table 4. Naturally, more false positives were introduced in the final reports when analyzing the projects in the testing dataset, even though the true positives remained the same. This in turn decreased the average F1 score to 0.29.
5.5.3. Generalization
Query generalization is important in the design of MoCQ. It enables MoCQ to capture diverse vulnerability patterns beyond the given input vulnerability example. Many new patterns could only be discovered with generalization. When we disabled it from MoCQ ($W/O_{Gen}$), the F1 score significantly dropped to 0.29. Our manual analysis of some queries revealed many new vulnerability patterns that the full-fledged MoCQ previously identified now disappeared. It is reasonable because by explicitly instrumenting the LLM to generalize the patterns, the LLM can better capture structural variations across vulnerabilities.
5.5.4. Models
MoCQ is not limited to only Claude 3.7 Sonnet. We further experimented with another model, GPT-4o (OpenAI, 2024). The results on Joern are shown in Table 4. In our experiments, GPT-4o has an average F1 score of 0.37, while full-fledged MoCQ with Claude 3.7 Sonnet has the best performance of 0.49. Our experience suggests that the Claude 3.7 Sonnet model has a better capability in coding tasks than GPT-4o. Though we did not integrate other models due to our cost constraints, we believe MoCQ has wide applicability.
5.5.5. Static Tools
Besides Joern, we experimented MoCQ with CodeQL. CodeQL is not capable of analyzing PHP. We thus evaluated it using the same setup for other vulnerabilities in our dataset. From our experience, CodeQL queries are often considered harder due to their higher DSL complexity, so the results on two types of vulnerabilities were worse than MoCQ with Joern. It is interesting to observe that MoCQ with CodeQL achieved a better F1 score (0.72) on JavaScript prototype pollution.
Table 4. Ablation study with F1 scores, experimented using 10 input projects. The default setting is with Joern and Claude 3.7 Sonnet.
Vul. Type MoCQ W/OFP W/OGen GPT-4o CodeQL DSL100
PHP SQLi 0.65 0.54 0.37 0.44 - 0.38
PHP XSS 0.52 0.32 0.25 0.43 - 0.23
PHP type. 0.25 0.19 0.23 0.31 - 0.15
PHP deser. 0.59 0.34 0.28 0.50 - 0.41
JS proto. 0.69 0.33 0.28 0.54 0.72 0.54
JS Cmd. 0.37 0.18 0.32 0.22 0.35 0.29
JS XSS 0.48 0.23 0.32 0.35 0.38 0.33
C/C++ UAF 0.53 0.31 0.35 0.38 0.44 0.28
C/C++ heap ov. 0.46 0.27 0.22 0.30 0.39 0.23
Java dan. func 0.50 0.29 0.33 0.35 0.48 0.36
Java SQLi 0.55 0.26 0.29 0.32 0.40 0.30
Java API misuse 0.50 0.25 0.28 0.31 0.36 0.22
Average 0.51 0.29 0.29 0.37 0.44 0.31
5.5.6. DSL Subset
Joern exposes a total of 4,5914,5914 , 591 DSL APIs, while CodeQL exposes 12,93412,93412 , 934 DSL APIs. The subsetting process prunes 48% of Joern’s and 67% of CodeQL’s APIs, substantially reducing the search space for query generation while maintaining sufficient expressiveness for effective vulnerability detection.
The number of input queries influences the resulting DSL subset. We provided Claude 3.7 Sonnet with increasing numbers of query examples and observed that the subset size stabilized around 12 to 16 queries for both Joern and CodeQL. We omit the detailed diagram showing the stabilization curve due to space constraints. By sampling different sets but the same number of queries, we found a high overlap (93.8%) in the resulting DSL subsets, indicating robustness against variability in query examples.
We further manually analyzed 50 randomly sampled removed APIs; all of them were confirmed redundant, showing the reliability of LLMs in the DSL subsetting task. Most (45 out of 50) removed features had equivalent alternatives within the retained subset. For example, Joern’s whereNot() can be expressed as where(not). A few removed APIs were non-essential features such as debugging or printing operations, and some were language-specific APIs irrelevant to the target languages (e.g., address used only in binary analysis).
To understand the effect of the subsetting technique, we developed a variant containing the full set of DSL, denoted as DSL100, and used it to generate queries. Generally, DSL100 resulted in much longer query generation times and frequently reached the attempt threshold before producing valid queries. This led to lower-quality merged queries and worse detection results, as shown in Table 4. We also experimented with removing an additional 10% of random APIs from the subset and found that many previously successful queries could no longer be generated. This highlights the importance of practical DSL subsetting to retain core, necessary features.
6. Discussion
Vulnerability Examples. We provided MoCQ with a small set of vulnerability examples to help it generate patterns for general vulnerability types. These examples serve as seed inputs from which MoCQ induces generalized patterns applicable to a broader class of similar vulnerabilities. They can be sourced from real-world data, such as CVE entries, security patches, or well-established vulnerability benchmarks. Importantly, the number of examples required is significantly smaller than those typically needed by deep learning-based techniques. Moreover, thanks to the code generation capabilities of modern LLMs, it is also feasible to generate high-quality synthetic vulnerability examples.
Extensibility. By design, MoCQ is language-agnostic and applicable to a wide range of languages and vulnerability types beyond those evaluated. It requires only a small number of vulnerability examples without requiring any changes to its design. Beyond Joern (Yamaguchi et al., 2014; Backes et al., 2017) and CodeQL (CodeQL, 2024a), MoCQ can be directly extended to frameworks such as Amazon CodeGuru (Services, 2025) and Datalog-based systems (Scholz et al., 2016). For other non-query-based static approaches, it may require specifying how the patterns are expressed and represented in these methods.
Other Use Scenarios. As a generic framework, MoCQ supports multiple practical use cases. For instance, it can enhance expert-developed queries by automatically extending them with additional detection logic, as demonstrated for JavaScript prototype pollution. MoCQ can also serve as a foundation for building a digital twin of closed-source static analyzers. In this setting, feedback such as true and false positives or partial matches can guide MoCQ to approximate the tool’s behavior. This is especially valuable in industrial contexts, where high-performing tools often lack transparency or extensibility. By mimicking such tools, MoCQ enables flexible, explainable analysis suitable for modern development workflows.
Ethical Considerations. We adhere to the best research ethics practices. We used LLMs in a local environment to generate vulnerability detection patterns. All newly discovered vulnerabilities and patterns were responsibly disclosed to the respective developers.
Threat to Validity. Our evaluation faces several threats to validity. First, our dataset is limited to 12 vulnerability types across four programming languages, and we did not experiment with others, though theoretically MoCQ is language-agnostic. Second, our comparison to expert-developed queries uses commit history as a proxy for development effort, which may not reflect actual time investment. Third, the manual collection of ground truth vulnerabilities may be incomplete, potentially affecting precision calculations.
7. Related Work
Learning-based Detection. Transformer-based models have been widely applied to vulnerability detection, including encoder-only (Fen, 2020), encoder-decoder (Wang et al., 2021, 2023a), and decoder-only (Touvron et al., 2023) architectures. To improve performance, researchers have explored various fine-tuning strategies, such as domain-specific pretraining (Ding et al., 2023), instruct-tuning (Mao et al., 2024; Chang et al., 2024), supervised fine-tuning (Zhu et al., 2024; Wang et al., 2024b), and parameter-efficient tuning (Mao et al., 2024). However, recent studies (Ullah et al., 2024) show that they require a large dataset for training and often present reduced effectiveness on under-represented vulnerabilities and large codebases. Compared to these approaches, MoCQ benefits from symbolic reasoning and achieves superior scalability, and requires only a small set of vulnerability examples.
Neuro-symbolic Program Analysis. Neuro-symbolic program analysis integrates neural and symbolic components. IRIS (Li et al., 2025) and Artemis (Ji et al., 2025) use LLMs to infer project-specific taint specifications for Java and PHP, respectively, based on signatures and comments. IRIS additionally applies LLMs to reduce false positives via slicing. Other works use neural components to assist symbolic validation, such as WAP (Medeiros et al., 2011), or decompose analysis into smaller property checks as in LLMSA (Wang et al., 2024a). LLift (Li et al., 2024) applies LLMs to reason about post-conditions in use-before-initialization bugs in the Linux kernel. MoCQ differs by aiming to automatically generate vulnerability queries from scratch, and can complement existing approaches by reducing manual engineering effort.
General Static Analysis. Static vulnerability detection includes a broad range of techniques beyond pattern-based approaches. Methods such as symbolic execution (Baldoni et al., 2018; Dudina and Belevantsev, 2017), model checking (Yang et al., 2006; Su et al., 2021), and formal verification (Hasan and Tahar, 2015) analyze program behavior against formal specifications or mathematical models, aiming to logically prove security properties or identify bugs. Other general techniques, such as dataflow analysis frameworks based on lattices (Cousot and Cousot, 1977), apply abstract interpretation to systematically reason about program behavior and detect potential vulnerabilities. Unlike pattern-based tools, these approaches do not depend on manually crafted vulnerability signatures. However, despite their precision and theoretical rigor, formal and dataflow-based methods often face significant scalability challenges when applied to large or complex codebases.
8. Conclusion
This work presents MoCQ, a novel neuro-symbolic approach that combines the complementary strengths of LLMs and static analysis to enable scalable, automated vulnerability detection. By leveraging a DSL subsetting technique and trace-driven symbolic validation, MoCQ automatically extracts vulnerability patterns using an LLM. MoCQ-generated queries achieved comparable detection performance to expert-developed queries with substantially reduced engineering effort. MoCQ lowers the barrier to developing custom vulnerability detectors and makes automated security analysis more accessible and adaptable. We believe this neuro-symbolic approach represents a promising direction for general software engineering.